# Debiased Inference for Dynamic Nonlinear Panels with Multi-dimensional Heterogeneities

### v3

@@ A.1 Lemmas
Let $\widetilde{u}_{i}^{\breve{\alpha}}(\theta)$ denote the $i$-th element of $H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)$ for $i=1,\ldots,N-1.$ Using (A.7) and (A.8), we can show that
$C_{NT,1}(\theta)$
$=$ $\frac{1}{N}[\widetilde{u}_{i}^{\breve{\alpha}}(\theta)]^{2}\bigg{(}\frac{1}{T}\sum_{t=1}^{T}\partial_{\alpha_{i}\alpha_{i}\alpha_{i}}l_{i,t}(\theta,\phi)|_{\phi=\phi(\theta)}\bigg{)}-\frac{1}{N}\bigg{(}\sum_{i=1}^{N-1}\widetilde{u}_{i}^{\breve{\alpha}}(\theta)\bigg{)}^{2}\bigg{(}\frac{1}{T}\sum_{t=1}^{T}\partial_{\alpha_{N}\alpha_{N}\alpha_{N}}l_{N,t}(\theta,\phi)|_{\phi=\phi(\theta)}\bigg{)}$
$=$ $O_{\mathbb{P}}\big{(}1/(NT)\big{)},$
uniformly in $\theta.$ By similar arguments, we also show that
$\sup_{\theta\in\Theta}C_{NT,2}(\theta)=O_{\mathbb{P}}\big{(}1/(NT)^{3/2}\big{)},\quad\sup_{\theta\in\Theta}C_{NT,3}(\theta)=O_{\mathbb{P}}\big{(}1/(NT)^{3/2}\big{)}.$
@@
$\sup_{\theta\in\Theta}\sup_{a_{j}\in(0,1)}r_{\psi_{j}}^{s}(\theta,a_{j})$
$=$ $-\frac{1}{2N}\bigg{(}\sum_{i=1}^{N-1}\widetilde{u}_{i}^{\breve{\alpha}}(\theta)\bigg{)}^{2}\bigg{(}\frac{1}{T}\sum_{t=1}^{T}\partial_{\alpha_{N}\alpha_{N}\alpha_{N}}l_{N,t}(\theta,\phi)|_{\phi=\phi(\theta)}\bigg{)}+o_{\mathbb{P}}\big{(}1/(NT)\big{)}$ (A.13)
$=$ $O_{\mathbb{P}}\big{(}1/(NT)\big{)}.$
Similarly, when $\psi_{j}=\gamma_{t},$ we have
$\sup_{\theta\in\Theta}\sup_{a_{j}\in(0,1)}r_{\psi_{j}}^{s}(\theta,a_{j})$
$=$ $-\frac{1}{2T}\bigg{(}\sum_{t=1}^{T-1}\widetilde{u}_{t}^{\breve{\gamma}}(\theta)\bigg{)}^{2}\bigg{(}\frac{1}{N}\sum_{i=1}^{N}\partial_{\gamma_{T}\gamma_{T}\gamma_{T}}l_{i,T}(\theta,\phi)|_{\phi=\phi(\theta)}\bigg{)}+o_{\mathbb{P}}\big{(}1/(NT)\big{)}$ (A.14)
$=$ $O_{\mathbb{P}}\big{(}1/(NT)\big{)},$
where $\widetilde{u}_{t}^{\breve{\gamma}}(\theta)$ denotes the $t$-th element of $H_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)$ for $t=1,\ldots,T-1.$ This shows the desired result.
@@
Under Assumption 1, $r(\theta)$ defined by (2.7) satisfies
@@
Proof. In the proof, we retain the notation and setup from Lemmas 1 and A.1. From Lemma A.1, we know that $NT\cdot r^{s}\left(\theta,a\right)$ is comparable to $(NT)^{3/4}\cdot s(\theta).$ Following similar arguments as in Lemma 1, we can show that
$\mathbb{E}\big{[}(r^{s}(\theta,a))^{\prime}[{H}\left(\theta\right)]^{-1}r^{s}(\theta,a)\big{]}=O\big{(}1/(NT)\big{)}.$
@@
$\big{(}\widehat{\psi}(\theta)-\psi(\theta)\big{)}^{\prime}r^{s}(\theta,a_{0}\,\iota_{N+T-2})$
$=$ $\frac{1}{2N}\bigg{(}\sum_{i=1}^{N-1}\widetilde{u}_{i}^{\breve{\alpha}}(\theta)\bigg{)}^{2}\bigg{(}\frac{1}{T}\sum_{t=1}^{T}\partial_{\alpha_{N}\alpha_{N}\alpha_{N}}l_{N,t}(\theta,\phi)|_{\phi=\phi(\theta)}\bigg{)}$
$+\frac{1}{2T}\bigg{(}\sum_{t=1}^{T-1}\widetilde{u}_{t}^{\breve{\gamma}}(\theta)\bigg{)}^{2}\bigg{(}\frac{1}{N}\sum_{i=1}^{N}\partial_{\gamma_{T}\gamma_{T}\gamma_{T}}l_{i,T}(\theta,\phi)|_{\phi=\phi(\theta)}\bigg{)}$
$\quad+\bigg{(}\sum_{i=1}^{N-1}\widetilde{u}_{i}^{\breve{\alpha}}(\theta)+\sum_{t=1}^{T-1}\widetilde{u}_{t}^{\breve{\gamma}}(\theta)\bigg{)}\cdot o_{\mathbb{P}}\big{(}1/(NT)\big{)}$
$=$ $O_{\mathbb{P}}\big{(}1/(NT)^{5/4}\big{)}.$ (A.15)
By Assumption 1 and applying the $L_{1}$ Convergence Theorem to (A.15), we find that
$\sup_{\theta\in\Theta}\mathbb{E}\big{[}\big{(}\widehat{\psi}(\theta)-\psi(\theta)\big{)}^{\prime}r^{s}\left(\theta,a_{0}\,\iota_{N+T-2}\right)\big{]}=o\big{(}1/(NT)\big{)}.$
@@
$\mathbb{E}\big{[}r(\theta)\big{]}=\frac{1}{2}\mathbb{E}\big{[}(r^{s}\left(\theta,a\right))^{\prime}[{H}(\theta)]^{-1}r^{s}\left(\theta,a\right)\big{]}+\frac{1}{3}\mathbb{E}\big{[}\big{(}\widehat{\psi}(\theta)-\psi(\theta)\big{)}^{\prime}r^{s}\left(\theta,a_{0}\,\iota_{N+T-2}\right)\big{]}=O\big{(}1/(NT)\big{)}$
@@
A.2 Proof of Theorem 1
From (2.7), we have
$\mathbb{E}[l(\theta)]=\mathbb{E}[\widehat{l}(\theta)]+\frac{1}{2}\mathbb{E}\big{\{}s^{\prime}(\theta)[H(\theta)]^{-1}s(\theta)\big{\}}-\mathbb{E}[r(\theta)].$
Thus, the desired result follows directly from Lemmas 1 and A.2.
A.3 Proof of Theorem 2
We retain the notation introduced in Lemmas 1-A.2. First, we show that
$\widehat{B}_{\breve{\alpha}}(\theta)-B_{\breve{\alpha}}(\theta)=o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)},\quad\text{and}\quad\widehat{B}_{\breve{\gamma}}(\theta)-B_{\breve{\gamma}}(\theta)=o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)}.$ (A.16)
Note that
$\widehat{B}_{\breve{\alpha}}(\theta)-B_{\breve{\alpha}}(\theta)=\frac{1}{2}\text{trace}\big{\{}\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}\left(\theta\right)[\Sigma^{\breve{\alpha}}(\theta)-\widehat{\Sigma}^{\breve{\alpha}}_{\tau}(\theta)]\big{\}}+\frac{1}{2}\text{trace}\big{\{}[\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)-H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)]\Sigma^{\breve{\alpha}}(\theta)\big{\}}.$
Define $\widetilde{s}_{i,t}^{\alpha}(\theta):=\{\partial_{\alpha_{i}}l_{i,t}(\theta,{\phi})-\mathbb{E}[\partial_{\alpha_{i}}l_{i,t}(\theta,{\phi})]\}|_{\phi={\phi}(\theta)}.$ Then we have
$N^{2}\big{[}{\Sigma}_{1}^{\breve{\alpha}}(\theta)-\widehat{\Sigma}_{1,\tau}^{\breve{\alpha}}(\theta)\big{]}$
@@ A.3 Proof of Theorem 2
[…]
$+\frac{1}{T^{2}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathds{1}\left\{\left|t-s\right|>\tau\right\}$
×$\qquad\times\text{diag}\big{\{}\mathbb{E}\big{[}\widetilde{s}_{1,t}^{\alpha}(\theta)\big{(}\widetilde{s}_{1,s}^{\alpha}(\theta)\big{)}^{\prime}\big{]},\ldots,\mathbb{E}\big{[}\widetilde{s}_{N-1,t}^{\alpha}(\theta)\big{(}\widetilde{s}_{N-1,s}^{\alpha}(\theta)\big{)}^{\prime}\big{]}\big{\}}.$ (A.17)
@@
$\mathbb{E}\big{[}\widetilde{s}_{i,t}^{\alpha}(\theta)\big{(}\widetilde{s}_{i,s}^{\alpha}(\theta)\big{)}^{\prime}\big{]}=\mathbb{E}\big{[}\widetilde{s}_{i,t}^{\alpha}(\theta)\widetilde{s}_{i,s}^{\alpha}(\theta)\big{]}\leq 8[a_{i}(|t-s|)]^{1-2/\eta}\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta},$
@@
$\frac{1}{T}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathds{1}\left\{\left|t-s\right|>\tau\right\}\mathbb{E}\big{[}\widetilde{s}_{i,t}^{\alpha}(\theta)\big{(}\widetilde{s}_{i,s}^{\alpha}(\theta)\big{)}^{\prime}\big{]}$
$\leq$ $8\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\cdot\frac{1}{T}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathds{1}\left\{\left|t-s\right|>\tau\right\}[a_{i}(|t-s|)]^{1-2/\eta}$
$\leq$ $8\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\cdot\frac{T-\tau}{T}\sum_{m=\tau+1}^{\infty}[a_{i}(m)]^{1-2/\eta}$
@@
given that $\tau/T\to 0.$ Combining this with (A.17) implies
$\sup_{\theta\in\Theta}N^{2}\|{\Sigma}_{1}^{\breve{\alpha}}(\theta)-\widehat{\Sigma}_{1,\tau}^{\breve{\alpha}}(\theta)\|_{\max}=o_{\mathbb{P}}(1/T).$ (A.18)
From the proof of Lemma A.1, we also have $\sup_{\theta\in\Theta}\|\widehat{\psi}(\theta)-\psi(\theta)\|_{\max}=o_{\mathbb{P}}(1).$ Moreover, under the conditional strong mixing condition, each element of $T^{-1}\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)$ converges in probability to its counterpart in $T^{-1}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)$ by the law of large numbers. Therefore, from (A.18),
$\text{trace}\big{\{}\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)[\Sigma_{1}^{\breve{\alpha}}(\theta)-\widehat{\Sigma}_{1,\tau}^{\breve{\alpha}}(\theta)]\big{\}}$
$=$ $\frac{T}{N^{2}}\text{trace}\big{\{}T^{-1}\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\,N^{2}[\Sigma^{\breve{\alpha}}_{1}(\theta)-\widehat{\Sigma}^{\breve{\alpha}}_{1,\tau}(\theta)]\big{\}}$
$=$ $\frac{T}{N}\cdot\frac{1}{N}\sum_{i=1}^{N-1}h_{i,i}^{\ast\breve{\alpha}}(\theta)\cdot o_{\mathbb{P}}(1/T)$
@@
Similarly,
$\text{trace}\big{\{}\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)[\Sigma_{2}^{\breve{\alpha}}(\theta)-\widehat{\Sigma}_{2,\tau}^{\breve{\alpha}}(\theta)]\big{\}}$
$=$ $\frac{T}{N^{2}}\text{trace}\big{\{}T^{-1}\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\,N^{2}[\Sigma_{2}^{\breve{\alpha}}(\theta)-\widehat{\Sigma}_{2,\tau}^{\breve{\alpha}}(\theta)]\big{\}}$
$=$ $\frac{T}{N^{2}}\cdot\sum_{i=1}^{N-1}\sum_{j=1}^{N-1}h_{i,j}^{\ast\breve{\alpha}}(\theta)\cdot o_{\mathbb{P}}(1/T)$
@@
where the last equality follows from the fact that $\sum_{j=1}^{N-1}h_{i,j}^{\ast\breve{\alpha}}(\theta)=O(1/N)$ uniformly in both $i$ and $\theta,$ as shown in Lemma A.1. Therefore,
$\text{trace}\big{\{}\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)[\Sigma^{\breve{\alpha}}(\theta)-\widehat{\Sigma}^{\breve{\alpha}}_{\tau}(\theta)]\big{\}}=o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)}.$ (A.19)
Moreover, from Theorem 1, we also have
$\text{trace}\big{\{}[\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)-H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)]\Sigma^{\breve{\alpha}}(\theta)\big{\}}=o_{\mathbb{P}}(1)\cdot\text{trace}\big{\{}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\Sigma^{\breve{\alpha}}(\theta)\big{\}}=o_{\mathbb{P}}(1)\cdot B_{\breve{\alpha}}(\theta)=o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)}.$ (A.20)
Combining (A.19) and (A.20) shows the first part of (A.16). The second part of (A.16) can be shown similarly.
@@
$l(\theta)=\widehat{l}(\theta)+\frac{1}{2}s_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)s_{\breve{\alpha}}(\theta)+\frac{1}{2}s_{\breve{\gamma}}^{\prime}(\theta)H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)s_{\breve{\gamma}}(\theta)+o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)}$

### v4

@@ A Proofs
Let $v_{i}^{\alpha}(\theta)$ denote the $i$-th element of $H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)$, for $i=1,\ldots,N-1$. Using (A.8) and (A.9), it can be shown that
$\quad\quad C_{NT,1}(\theta)$
$=\frac{1}{N}\left(\sum_{i=1}^{N-1}v_{i}^{\alpha}(\theta)\right)^{2}\left(\frac{1}{T}\sum_{t=1}^{T}\partial_{\alpha_{i}\alpha_{i}\alpha_{i}}l_{i,t}(\theta,\phi)|_{\phi=\phi(\theta)}\right)-\frac{1}{N}\left(\sum_{i=1}^{N-1}v_{i}^{\alpha}(\theta)\right)^{2}\left(\frac{1}{T}\sum_{t=1}^{T}\partial_{\alpha_{N}\alpha_{N}\alpha_{N}}l_{N,t}(\theta,\phi)|_{\phi=\phi(\theta)}\right)$
$=O_{\mathbb{P}}\left(1/(NT)\right),$
uniformly in $\theta$. By similar arguments, we also have
$\sup_{\theta\in\Theta}C_{NT,2}(\theta)=O_{\mathbb{P}}\left(1/(NT)^{3/2}\right),\qquad\sup_{\theta\in\Theta}C_{NT,3}(\theta)=O_{\mathbb{P}}\left(1/(NT)^{3/2}\right).$
@@
$\quad\quad\sup_{\theta\in\Theta}\sup_{a_{j}\in(0,1)}r_{\psi_{j}}^{s}(\theta,a_{j})$
$=-\frac{1}{2N}\left(\sum_{i=1}^{N-1}v_{i}^{\alpha}(\theta)\right)^{2}\left(\frac{1}{T}\sum_{t=1}^{T}\partial_{\alpha_{N}\alpha_{N}\alpha_{N}}l_{N,t}(\theta,\phi)|_{\phi=\phi(\theta)}\right)+o_{\mathbb{P}}\left(1/(NT)\right)$
$=O_{\mathbb{P}}\left(1/(NT)\right).$ (A.14)
Similarly, when $\psi_{j}=\gamma_{t}$ ($t=1,\ldots,T-1$),
$\quad\quad\sup_{\theta\in\Theta}\sup_{a_{j}\in(0,1)}r_{\psi_{j}}^{s}(\theta,a_{j})$
$=-\frac{1}{2T}\left(\sum_{t=1}^{T-1}v_{t}^{\gamma}(\theta)\right)^{2}\left(\frac{1}{N}\sum_{i=1}^{N}\partial_{\gamma_{T}\gamma_{T}\gamma_{T}}l_{i,T}(\theta,\phi)|_{\phi=\phi(\theta)}\right)+o_{\mathbb{P}}\left(1/(NT)\right)$
$=O_{\mathbb{P}}\left(1/(NT)\right),$ (A.15)
where $v_{t}^{\gamma}(\theta)$ denotes the $t$-th element of $H_{\gamma\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)$ for $t=1,\ldots,T-1$. This leads to the desired result.
The next lemma states that the final remainder, $r(\theta)$ in Equation (2.14), is $O\left(1/\left(NT\right)\right)$ uniformly in $\theta$, where
$r(\theta)$ $:=\frac{1}{2}(r^{s}(\theta,a))^{\prime}[H\left(\theta\right)]^{-1}r^{s}(\theta,a)+r^{l}(\theta,a_{0}),$ (A.16)
$r^{l}\left(\theta,a_{0}\right)$ $:=\frac{1}{3}\left(\widehat{\psi}(\theta)-\psi(\theta)\right)^{\prime}r^{s}(\theta,a_{0}\times\iota_{(N+T-2)K})$ (A.17)
for some $a_{0}\in(0,1)$. Here $r^{l}\left(\theta,a_{0}\right)=:r^{l}(\theta)$ as in Equation (2.12) from the main text. The dependency of $r(\theta)$ on $a$ and $a_{0}$ is suppress for simplicity here.
@@
Under Assumption 1, $r(\theta)$ defined by (A.16) satisfies
@@
Proof. In the proof, we retain the notation and setup from Lemmas A.1 and A.2. From Lemma A.2, we know that $NT\cdot r^{s}\left(\theta,a\right)$ is comparable to $(NT)^{3/4}\cdot s(\theta)$. Following similar arguments as in Lemma A.1, we can show that
$\mathbb{E}\left\{(r^{s}(\theta,a))^{\prime}[{H}\left(\theta\right)]^{-1}r^{s}(\theta,a)\right\}=O\left(1/(NT)\right).$
@@
$\quad\quad\left(\widehat{\psi}(\theta)-\psi(\theta)\right)^{\prime}r^{s}(\theta,a_{0}\,\iota_{N+T-2})$
$=\frac{1}{2N}\left(\sum_{i=1}^{N-1}v_{i}^{\alpha}(\theta)\right)^{2}\left(\frac{1}{T}\sum_{t=1}^{T}\partial_{\alpha_{N}\alpha_{N}\alpha_{N}}l_{N,t}(\theta,\phi)|_{\phi=\phi(\theta)}\right)$
$\quad\quad+\frac{1}{2T}\left(\sum_{t=1}^{T-1}v_{t}^{\gamma}(\theta)\right)^{2}\left(\frac{1}{N}\sum_{i=1}^{N}\partial_{\gamma_{T}\gamma_{T}\gamma_{T}}l_{i,T}(\theta,\phi)|_{\phi=\phi(\theta)}\right)$
$\quad\quad+\left(\sum_{i=1}^{N-1}v_{i}^{\alpha}(\theta)+\sum_{t=1}^{T-1}v_{t}^{\gamma}(\theta)\right)\cdot o_{\mathbb{P}}\left(1/(NT)\right)$
$=O_{\mathbb{P}}\left(1/(NT)^{5/4}\right).$ (A.18)
By Assumption 1 and applying the $L_{1}$ convergence theorem to A.18, we find that
$\sup_{\theta\in\Theta}\mathbb{E}\left[\left(\widehat{\psi}(\theta)-\psi(\theta)\right)^{\prime}r^{s}\left(\theta,a_{0}\,\iota_{N+T-2}\right)\right]=o\left(1/(NT)\right).$
@@
$\mathbb{E}\left[r(\theta)\right]=\frac{1}{2}\mathbb{E}\left[(r^{s}\left(\theta,a\right))^{\prime}[{H}(\theta)]^{-1}r^{s}\left(\theta,a\right)\right]+\frac{1}{3}\mathbb{E}\left[\left(\widehat{\psi}(\theta)-\psi(\theta)\right)^{\prime}r^{s}\left(\theta,a_{0}\,\iota_{N+T-2}\right)\right]=O\left(1/(NT)\right)$
@@
Next, we prove Theorem 1 and 2 below.
Proof of Theorem 1. Following the expansion in (2.12) and (2.13), we have
$\mathbb{E[}\widehat{l}(\theta)-l(\theta)]=-\frac{1}{2}\mathbb{E\{}s^{\prime}(\theta)[H(\theta)]^{-1}s(\theta)\}+\mathbb{E}r(\theta)$
for $r\left(\theta\right)$ defined in (A.16). The claim in Theorem 1 follows directly from Lemmas A.1 and A.3.
The proof of Theorem 2 is as follows.
Proof of Theorem 2. We retain the notation introduced in Lemmas A.1-A.3. We first show
$\widehat{B}_{\alpha}(\theta)-B_{\alpha}(\theta)=o_{\mathbb{P}}(1/\sqrt{NT}),\qquad\widehat{B}_{\gamma}(\theta)-B_{\gamma}(\theta)=o_{\mathbb{P}}(1/\sqrt{NT}).$ (A.19)
To show the first part of (A.19), note that
$\widehat{B}_{\alpha}(\theta)-B_{\alpha}(\theta)=\frac{1}{2}\operatorname*{trace}\left\{\widehat{H}_{\alpha\alpha^{\prime}}^{\ast}\left(\theta\right)\left[S_{\alpha\alpha}\left(\theta\right)-\widehat{S}_{\alpha\alpha}\left(\theta\right)\right]\right\}+\frac{1}{2}\operatorname*{trace}\left\{\left[\widehat{H}_{\alpha\alpha^{\prime}}^{\ast}(\theta)-H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\right]S_{\alpha\alpha}\left(\theta\right)\right\},$
where
$S_{\alpha\alpha}\left(\theta\right)$ $=S_{\alpha\alpha}^{1}\left(\theta\right)+S_{\alpha\alpha}^{2}\left(\theta\right),$
$S_{\alpha\alpha}^{1}\left(\theta\right)$ $:=\operatorname*{diag}\left\{\mathbb{E}\left[s_{1}^{\alpha}\left(\theta\right)\left(s_{1}^{\alpha}\left(\theta\right)\right)^{\prime}\right],\ldots,\mathbb{E}\left[s_{N-1}^{\alpha}\left(\theta\right)\left(s_{N-1}^{\alpha}\left(\theta\right)\right)^{\prime}\right]\right\},$
$S_{\alpha\alpha}^{2}\left(\theta\right)$ $:=(\iota_{N-1}\iota_{N-1}^{\prime})\otimes\mathbb{E}\left[s_{N}^{\alpha}\left(\theta\right)\left(s_{N}^{\alpha}\left(\theta\right)\right)^{\prime}\right];$
$\widehat{S}_{\alpha\alpha}\left(\theta\right)$ $=\widehat{S}_{\alpha\alpha}^{1}\left(\theta\right)+\widehat{S}_{\alpha\alpha}^{2}\left(\theta\right),$
$S_{\alpha\alpha}^{1}\left(\theta\right)$ $:=\operatorname*{diag}\left\{\widehat{s}_{1}^{\alpha}\left(\theta\right)\left(\widehat{s}_{1}^{\alpha}\left(\theta\right)\right)^{\prime},\ldots,\widehat{s}_{N-1}^{\alpha}\left(\theta\right)\left(\widehat{s}_{N-1}^{\alpha}\left(\theta\right)\right)^{\prime}\right\},$
(${}_{\alpha\alpha}^{2}\left(\theta\right)$ $:=(\iota_{N-1}\iota_{N-1}^{\prime})\otimes\left[\widehat{s}_{N}^{\alpha}\left(\theta\right)\left(\widehat{s}_{N}^{\alpha}\left(\theta\right)\right)^{\prime}\right].$
Here we have
$\quad\quad N^{2}\left[S_{\alpha\alpha}^{1}\left(\theta\right)-\widehat{S}_{\alpha\alpha}^{1}\left(\theta\right)\right]$
@@
×$\quad\quad\times\operatorname*{diag}\left\{\widetilde{s}_{1,t}^{\alpha}(\theta)\left(\widetilde{s}_{1,s}^{\alpha}(\theta)\right)^{\prime}-\widehat{s}_{1,t}^{\alpha}(\theta)\left(\widehat{s}_{1,s}^{\alpha}(\theta)\right)^{\prime},\ldots,\widetilde{s}_{N-1,t}^{\alpha}(\theta)\left(\widetilde{s}_{N-1,s}^{\alpha}(\theta)\right)^{\prime}-\widehat{s}_{N-1,t}^{\alpha}(\theta)\left(\widehat{s}_{N-1,s}^{\alpha}(\theta)\right)^{\prime}\right\}$
$\quad\quad+\frac{1}{T^{2}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathds{1}\left\{\left|t-s\right|>\tau\right\}$
×$\quad\quad\times\operatorname*{diag}\left\{\mathbb{E}\left[\widetilde{s}_{1,t}^{\alpha}(\theta)\left(\widetilde{s}_{1,s}^{\alpha}(\theta)\right)^{\prime}\right],\ldots,\mathbb{E}\left[\widetilde{s}_{N-1,t}^{\alpha}(\theta)\left(\widetilde{s}_{N-1,s}^{\alpha}(\theta)\right)^{\prime}\right]\right\}.$ (A.20)
@@
$\mathbb{E}\left[\widetilde{s}_{i,t}^{\alpha}(\theta)\left(\widetilde{s}_{i,s}^{\alpha}(\theta)\right)^{\prime}\right]\leq 8[a_{i}(|t-s|)]^{1-2/\eta}\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta},$
@@
$\quad\quad\frac{1}{T}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathds{1}\left\{\left|t-s\right|>\tau\right\}\mathbb{E}\left[\widetilde{s}_{i,t}^{\alpha}(\theta)\left(\widetilde{s}_{i,s}^{\alpha}(\theta)\right)^{\prime}\right]$
$\leq 8\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\left(\frac{1}{T}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathds{1}\left\{\left|t-s\right|>\tau\right\}[a_{i}(|t-s|)]^{1-2/\eta}\right)$
$\leq 8\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\left(\frac{T-\tau}{T}\sum_{m=\tau+1}^{\infty}[a_{i}(m)]^{1-2/\eta}\right)$
@@
given that $\tau/T\rightarrow 0$. Combining this with (A.20), we have
$\sup_{\theta\in\Theta}N^{2}\left\|S_{\alpha\alpha}^{1}\left(\theta\right)-\widehat{S}_{\alpha\alpha}^{1}\left(\theta\right)\right\|_{\max}=o_{\mathbb{P}}(1/T).$ (A.21)
From the proof of Lemma A.2, we also have $\sup_{\theta\in\Theta}\|\widehat{\psi}(\theta)-\psi(\theta)\|_{\max}=o_{\mathbb{P}}(1)$. Moreover, under the conditional strong mixing assumption, each element of $T^{-1}\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)$ converges in probability to the corresponding element in $T^{-1}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)$ by the law of large numbers. Therefore (A.21) further leads to
$\quad\quad\operatorname*{trace}\left\{\widehat{H}_{\alpha\alpha^{\prime}}^{\ast}\left(\theta\right)\left[S_{\alpha\alpha}^{1}\left(\theta\right)-\widehat{S}_{\alpha\alpha}^{1}\left(\theta\right)\right]\right\}$
$=\frac{T}{N^{2}}\operatorname*{trace}\left\{T^{-1}\widehat{H}_{\alpha\alpha^{\prime}}^{\ast}(\theta)\cdot N^{2}\left[S_{\alpha\alpha}^{1}\left(\theta\right)-\widehat{S}_{\alpha\alpha}^{1}\left(\theta\right)\right]\right\}$
$=\frac{T}{N}\frac{1}{N}\sum_{i=1}^{N-1}h_{i,i}^{\ast\alpha}(\theta)\cdot o_{\mathbb{P}}(1/T)$
@@
A similar procedure can be followed to show
$\quad\quad\operatorname*{trace}\left\{\widehat{H}_{\alpha\alpha^{\prime}}^{\ast}\left(\theta\right)\left[S_{\alpha\alpha}^{2}\left(\theta\right)-\widehat{S}_{\alpha\alpha}^{2}\left(\theta\right)\right]\right\}$
$=\frac{T}{N^{2}}\operatorname*{trace}\left\{T^{-1}\widehat{H}_{\alpha\alpha^{\prime}}^{\ast}(\theta)\cdot N^{2}\left[S_{\alpha\alpha}^{2}\left(\theta\right)-\widehat{S}_{\alpha\alpha}^{2}\left(\theta\right)\right]\right\}$
$=\frac{T}{N^{2}}\sum_{i=1}^{N-1}\sum_{j=1}^{N-1}h_{i,j}^{\ast\alpha}(\theta)\cdot o_{\mathbb{P}}(1/T)$
@@
where the last equality follows from, as shown in Lemma A.2, $\sum_{j=1}^{N-1}h_{i,j}^{\ast\alpha}(\theta)=O(1/N)$ uniformly in both $i$ and $\theta$. The above leads to
$\operatorname*{trace}\left\{\widehat{H}_{\alpha\alpha^{\prime}}^{\ast}(\theta)\left[S_{\alpha\alpha}\left(\theta\right)-\widehat{S}_{\alpha\alpha}\left(\theta\right)\right]\right\}=o_{\mathbb{P}}(1/\sqrt{NT}).$ (A.22)
Second, from Theorem 1, we also have
$\operatorname*{trace}\left\{\left[\widehat{H}_{\alpha\alpha^{\prime}}^{\ast}(\theta)-H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\right]S_{\alpha\alpha}\left(\theta\right)\right\}=o_{\mathbb{P}}(1)\cdot\operatorname*{trace}\left\{H_{\alpha\alpha^{\prime}}^{\ast}(\theta)S_{\alpha\alpha}\left(\theta\right)\right\}=o_{\mathbb{P}}(1)\cdot B_{\alpha}(\theta)=o_{\mathbb{P}}(1/\sqrt{NT}).$ (A.23)
Combining A.22 and A.23 leads to the first part of A.19. The second part of A.19 can be shown similarly.
@@
$l(\theta)=\widehat{l}(\theta)+\frac{1}{2}s_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)s_{\alpha}(\theta)+\frac{1}{2}s_{\gamma}^{\prime}(\theta)H_{\gamma\gamma^{\prime}}^{\ast}(\theta)s_{\gamma}(\theta)+o_{\mathbb{P}}(1/\sqrt{NT})$